The table below shows the number of male and female students enrolled in nursing at a universit Nursing majors Non-nursing majors Total Males Females Total 98 600 698 1019 1722 2741 2322 3439 (a) Find the probability that the student is male or a nursing major P(being male or being nursing major) (Round to the nearest thousandth as needed.) (b) Find the probability that the student is female or not a nursing major P(being female or not being a nursing major)- (Round to the nearest thousandth as needed.) (c) Find the probability that the student is not female or a nursing major Pnot being female or being a nursing major)- Round to the nearest thousandth as needed (d) Are the events being male and being a nursing major mutually exclusive? Explain. O A. No. because there are 98 males majoring in nursing O B. No because one cant be male and a nursing major at the same time O c. Yes, because one cant be male and a nursing major at the same time 0 D. Yes, because there are 98 males majoring in nursing
Answer #1

a) P(Male or Nursing major) = (1117 + 698 - 98) / 3439

= 0.4993

b) P(Female or Not a nursing major) = (2322 + 2421 - 1722) / 3439

= 0.9715

c) P(Not female or a nursing major) = (1117 + 698 - 98) / 3439

= 0.4993

d) Answer A : No, because there are 98 males majoring in nursing.
Answer #2

SOLUTION :


a.


P (Male or Nursing major) 

= P(Male) + P(nursing major) - P(Male and Nursing major)

= 1117/3439 + 698/3439 - 98/3439

= (1117+698 - 98)/3439

= 0.499 (ANSWER)


b.


P(Female or Non-nursing major)

= P(Female) + P(Non-nursing major) - P(Female and Non-nursing major)

= 2322/3439 + 2741/3439 - 1722/3439

= (2322 + 2741 - 1722)/3439

= 0.972 (ANSWER).


c.


P(Not-female or Nursing major)

= P(Not-female) + P(Nursing major) - P(Not-female and Nursing major)

= 1117/3439 + 698/3439 - 98/3439

= (1117+698 -m98)/3439

= 0.499 (ANSWER).


d.


A . No, because there are 98 males doing Nursing major. (ANSWER).


answered by: Tulsiram Garg